OpenStudy (anonymous):

Please help. Which is correct. Solve -2x2 +3x - 9 = 0 x equals quantity of 3 plus or minus 3i square root of 7 all over 4 x equals quantity of 3 plus or minus 9i square root of 7 all over 4 x equals quantity of negative 3 plus or minus 3i square root of 7 all over 4 x equals quantity of negative 3 plus or minus 9i square root of 7 all over 4 I will give medal :)

OpenStudy (anonymous):

@insa help?

OpenStudy (insa):

umm....i have no idea.. havent studied it yet :/

OpenStudy (anonymous):

Aww haha ok, thanks anyways @insa

OpenStudy (insa):

haha :D

OpenStudy (mathmale):

Please actually solve the equation -2x2 +3x - 9 = 0. You could make it easier for yourself by changing all of the signs: 2x2 - 3x + 9 = 0 How would you go about solving this? Hints: Some appropriate methods are 1. factoring 2. quadratic formula 3. graphing

OpenStudy (anonymous):

well the thing is, I have not really learned this yet? its a pre-study guide so could you help me? @mathmale

OpenStudy (mathmale):

Are you familiar with the quadratic formula? If so, type it out here.

OpenStudy (anonymous):

isn't it something like x= -b + sqrt b^2 -4ac/2a?

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

That's right! Good. Obtain your values of a, b and c from your original posting: -2x2 +3x - 9

OpenStudy (anonymous):

ummm x=-3 + sqrt 3^2 -4(-2)9/2a ?? lol Im sorry if its wrong 0.0 @mathmale

OpenStudy (mathmale):

Just compare: -2x2 +3x - 9 to ax^2 +bx + c and decide upon the values of a, b and c.

OpenStudy (mathmale):

Hint: c = -9. a = ? b = ??

OpenStudy (anonymous):

a= 2 and b=3? @mathmale

OpenStudy (anonymous):

*-2

OpenStudy (mathmale):

Yes: a=-2 b=3 c=-9 Now the quantity under the radical sign is called the "determinant," and the "determinant" equals b^2 - 4*a*c. Bet you've seen this before. Substitute the values of a, b and c in to this formula and tell me what you get.

OpenStudy (mathmale):

b^2, when b=3, is: (3)^2 = ?

OpenStudy (anonymous):

that would equal 9?

OpenStudy (mathmale):

Yes, and -4 ( a )*( c ) = ?

OpenStudy (anonymous):

a= 2 and c= -9 so -4(2)*(-9) -8*(-9) =-72? ._.

OpenStudy (mathmale):

Yes, and so your b^2 - 4ac has what value? Go back and take what you've already found for b^2 and -4ac.

OpenStudy (anonymous):

9-72? is that what you mean

OpenStudy (mathmale):

Indeed it is. 9-72= ?

OpenStudy (anonymous):

thats -63

OpenStudy (mathmale):

that's right. Can we find the square root of a negative number?

OpenStudy (anonymous):

7.93725393 i

OpenStudy (mathmale):

But that 63 is negative. Can we find the square root of a negative number?

OpenStudy (anonymous):

I'm guessing no?

OpenStudy (mathmale):

Are you familiar with the operator " i " found in situations such as this one? If you have Sqrt(-63), you re-write it as i*Sqrt(63). Have you seen that before?

OpenStudy (anonymous):

i believe so

OpenStudy (mathmale):

" i " is called the "imaginary operator."

OpenStudy (anonymous):

oh ok

OpenStudy (mathmale):

So, our Sqrt(-63) comes out to\[i*3*\sqrt{7}\] because sqrt(9)=3. Make sense to you?

OpenStudy (anonymous):

yes :)

OpenStudy (mathmale):

Great! Look at the four answer choices: x equals quantity of 3 plus or minus 3i square root of 7 all over 4 x equals quantity of 3 plus or minus 9i square root of 7 all over 4 x equals quantity of negative 3 plus or minus 3i square root of 7 all over 4 x equals quantity of negative 3 plus or minus 9i square root of 7 all over 4 We have to decide which one is correct. Note that your quadratic formula comes out to \[x=\frac{ -b. plus.or.minus.Sqrt(4ac) }{ 2a }\]and that you already know that Sqrt(4ac) = i*3*Sqrt(7).

OpenStudy (mathmale):

You know the values of b and c. So, would you write out your own expression for the roots, x, and then compare y our result to the four possible answers?

OpenStudy (anonymous):

i would say it would be the 2nd one?

OpenStudy (mathmale):

Have you actually written out your solution statement, x=(-b plus or minus ) Sqrt(b^2 - 4ac) ) / 2a ?

OpenStudy (anonymous):

wait, so it would be the first one then right?!

OpenStudy (mathmale):

Hold it, please...I know you want to finish as soon as possible, but I'd really appreciate your answering my question: have you actually written out your solution statement?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

and you think that the first of the four possible answers matches your result?

OpenStudy (anonymous):

yes i do

OpenStudy (mathmale):

but isn't b = 3, so that -b = -(3) = -3 ? Wouldn't that eliminate the first 2 answer choices?

OpenStudy (anonymous):

OH! i completely forgot the - I'm sorry

OpenStudy (mathmale):

In the end I'm going to let you make the choice of which answer is best. You've done very well here.

OpenStudy (anonymous):

I got it now but thank you so much! @mathmale :))))))

OpenStudy (mathmale):

We went through a lot of effort, but that's really the only way to learn this material. You're very welcome!!! Hope to see you again soon on OpenStudy. :)